OpenStudy (ttp):

A triangle has a leg that measures 6 and a hypotenuse of twelve. Which leg measure would form a right triangle? one hundred eight square root of one hundred eight one hundred eighty square root of one hundred eighty

OpenStudy (ttp):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

 a^2 + b^2 = c^2 6^2 + b^2 = 12^2 36 + b^2 = 144 -36 -36 b^2 = 108 That being the square root because you're solving for one of the sides.